Operations Report – January 4th

Crew 238 Operations Report 04-01-2022


Name of person filing report: Simon Werner

Non-nominal systems: 1 CO- Sensor, lower deck, replacement in progress
(Atila), spacesuits, radio, please see details below.

Notes on non-nominal systems: Sensor broken on 2022/01/03, spacesuits/
radio defects detected today at 6.30 pm during post EVA checks, see


Spirit rover used: Yes

Hours: (taken after Return from Marble Ritual): 149.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: yes

Hours: 79.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: yes

Hours: 158.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Perseverance rover used: no

Hours: unknown

Beginning charge: unknown

Ending charge: unknown

Currently charging: unknown

General notes and comments: Since today we are in SIM. A re-supply
mission from the outpost brought a new toilet. We’ve had an EVA to
Marble Ritual in the morning and afternoon, used Rovers, see EVA
report. Used rovers worked nominal, all plugged in. Rock samples have
been taken close to marble ritual, volvanic material (to be
estimated), Quartz, Flintstone and Achat (German).

Summary of Hab operations:


Water (static tank): 480

Water (loft tank): 28 gallons, filled to 55 gallons

Water Meter: post pumping 01544074

Static to Loft Pump used – yes

Static tank pipe heater (on or off): installation by Outpost, unknown status

Static tank heater (On or off) unknown

Toilet tank emptied: no

Summary of internet: nominal

Summary of suits and radios: Several spacesuit issues occurred.

Spacesuits next to airlock: Spacesuit nr. 2 is non-functional although
plugged in, will be checked in detail tomorrow. Spacesuit nr. 4 is on
low airflow, plugged in, will be re-checked tomorrow, Spacesuit nr. 5
was on low airflow, plugged to a different power socket, now full air
blow capacity availlable.

Spacesuits in the lower deck: Generation 1 type Spacesuit is on low
air flow, plugged in, re-check tomorrow, System nr. is out of order,
battery voltage reads + 5.1 Volts, fuse is ok (0.1 Ohms). Removed fuse
to check with the following connections: battery ground to multimeter
common, fuse into multimeter voltage (DC) in. Test procedure: switch
on, power supply plugged in. Voltage read was very instable, changing
from 0.0V to 13.75V. Repeatedly checked with 3 different power
supplies, same result. I suspect the power in socket or the switch to
have a half-lose cable connection. Each movement on the power supply
cable led to a different or even no voltage read. Suggested procedure:
move System to the RAM tomorrow, dis-assemble the Switch/plug unit and
check soldered connections/ for broken cables.

Same may be necessary for Spacesuit Nr. 1 and 4. Tomorrow is a day in
HAB, so here is time for this. All other spacesuits are nominal, incl.
those used during EVA.

Radio issues: One Midland-type radio (the yellow ones) does not
re-load, when taken out of the supply, it turns of. Asserted, tagged
as non-functional. Maybe the batteries died, re-supply of rechargeable
batteries, test again? All other radio systems are nominal

Summary of GreenHab operations:

WATER USE: 210 gallons (guess)

Heater: On

Supplemental light: On

Harvest: none

Summary of ScienceDome operations: none

Dual split: status unknown, operated by outpost

Summary of RAM operations: Took screwdriver set back to RAM, checked
Lights and ventilators. All nominal, except fan in the airlock. Makes
some heavy noise/ sounds when turned on –> mechanical contact of fan
blades to frame? Should this be checked in detail?

Ram tools have been used to make a military style/survival kitchen
grater from a can – lid. Noticed the following instrument is not
working: dremel tool, motor can be switched on, but no rotation at the
end of the flexible arm.

Summary of any observatory issues: Still missing the astronomy box,
see Operations report Sol1, Image of Messier galaxy taken by crew
astronomer, details to astronomy operations are to be found in
astronomy report.

Summary of health and safety issues: none, see HSO report

Questions, concerns and requests to Mission Support:

All re-supply delivered, thank you.



Copyright © The Mars Society. All rights reserved. | Main Site